Two cars X and Y accelerate at the rate of 2 m/s2 and 3 m/s2 respectively from rest.The ratio of time taken by the cars X and Y is 4:5.In that given ratio of time interval if the distance travelled by car x is 100 km then the distance travelled by car y is:

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT

Cars x velocity is 2 and y 3 ..

The ration of time travel is 4:5 ...

So ans is ...

Let take time travel by x is 4c. And travelling by y is 5c ..

Then ... V=d/t then d=vt so .. for x. .

4c×2=100 then c =100/8 ... C=25/2 ..

Then for y 5c ×3 =d .. so 125×3/2 =d .... So d = 375/2 ... Then d= 187.5 km 

THANKS